A lens, made of crown glass, has one convex surface with radius of curvature of 80.0 cm. What should be the radius of curvature of the other surface so that the focal length of the lens is 120 cm?
1
Expert's answer
2020-10-15T10:47:16-0400
"\frac{1}{f}=(n-1)\left(\frac{1}{R_1}+\frac{1}{R_2}\right)\\\frac{1}{120}=(1.5-1)\left(\frac{1}{80}+\frac{1}{R_2}\right)\\R_2=240\ cm"